OpenStudy (anonymous):

the ordered pairs (1,2) (2,4) (3,8) (4,16) (5,32) represents a function. What is the rule that represents this function?

OpenStudy (anonymous):

you need to look at the x and then look at the y, how are they changing?

OpenStudy (anonymous):

someone told me that there is suppose to be an exponent in there is it true thought because I tried exponents and then ended up getting confused

OpenStudy (anonymous):

yes there is an exponent but with a simple base

OpenStudy (anonymous):

Alright, so the question is simply asking us to find the function associated. Think of it this way, an in and out machine. you stick the x in and do something to it to get the next x value and the same for y. In general terms our change is represented by x or n, a number. So we see our x increases by 1, so we will leave that alone for now. If we look at our y, we see that it's not that easy. Each coordinate represents an nth pattern. here we have our first 5 pattern pieces. Stick that into 2^x so, for n=1, 2^1=2. n=2, 2^2=4, n=3, 2^3=8, n=4, 2^4=16, n=5, 2^5=32. So the correct formula is y=\[2^{x}\]

OpenStudy (anonymous):

thank you that makes soooo mmuch sense now lol I've been imputing it the wrong way